### v1

# Perturbative Distinguishability of Black Hole Microstates from AdS/CFT Correspondence

Abstract
We establish direct evidence for the perturbative distinguishability between black hole microstates and thermal states using the AdS/CFT correspondence. In two-dimensional holographic conformal field theories, we obtain the short interval expansion of subsystem fidelity and quantum Jensen-Shannon divergence, both of which provide rigorous lower and upper bounds for trace distance. This result demonstrates that quantum gravity corrections break semiclassical indistinguishability, thereby supporting the recovery of information even from a small amount of the Hawking radiation.
I Introduction
The black hole information paradox Hawking:1974sw ; Hawking:1976ra , the apparent conflict between unitary quantum evolution and thermal Hawking radiation, remains a cornerstone challenge in quantum gravity. While the AdS/CFT correspondence Maldacena:1997re ; Witten:1998qj ; Gubser:1998bc ; Aharony:1999ti provides a holographic framework for unitary evolution, translating this principle into operational distinguishability of microstates and thermal states requires confronting deep questions about semiclassical observables. Holographic entanglement entropy Ryu:2006bv ; Hubeny:2007xt ; Faulkner:2013ana and recent advances of the island formula Penington:2019npb ; Almheiri:2019psf have resolved the entropy paradox via the Page curve Page:1993df , yet fundamental gaps persist: How do quantum gravity corrections encode microscopic state differences into subsystems accessible to observers?
Within AdS3/CFT2, black hole microstates are dual to high-energy eigenstates in the boundary conformal field theory (CFT) Strominger:1997eq , while thermal ensembles describe equilibrium black holes with Bekenstein-Hawking entropy. Semiclassical arguments suggest local indistinguishability between these states, a manifestation of the eigenstate thermalization hypothesis (ETH) Deutsch:1991msp ; Srednicki:1994mfb ; Deutsch:2018thx ; Garrison:2015lva ; Dymarsky:2016ntg , where few-body observables in chaotic systems lose memory of microscopic details. This thermalization persists at leading order in holographic CFTs: highly excited primary states exhibit entanglement entropy matching thermal predictions at $O(c)$ Fitzpatrick:2014vua ; Fitzpatrick:2015zha , with the large central charge $c=\frac{3R}{2G}$ Brown:1986nw encoding the AdS3 radius $R$ and Newton constant $G$. The distinction becomes critical for small subsystems: while the Holevo information Bao:2017guc suggests classical indistinguishability at leading order, quantum corrections may encode state-specific signatures through $1/c$ corrections. Results of Rényi entropy Lashkari:2016vgj ; Lin:2016dxa , entanglement entropy and subsystem relative entropy He:2017vyf , subsystem Schatten distances Basu:2017kzo ; He:2017txy , and Holevo information Guo:2018djz indicate quantum gravity distinctions that are invisible to semiclassical probes.
Fixed area states Akers:2018fow ; Dong:2018seb ; Dong:2019piw , bulk geometries with constrained extremal surface areas, offer insights into state distinguishability. Studies reveal nonperturbative distinguishability scaling Kudler-Flam:2021rpr ; Kudler-Flam:2021alo , but their status as superpositions of eigenstates introduces ambiguities dependent on microscopic coefficients. In contrast, primary states are exact energy eigenstates, providing a natural arena to test intrinsic microstate properties.
The subsystem distinguishability of black hole microstates faces three unresolved challenges. Prior analyses focused on entropy differences Lashkari:2016vgj ; Lin:2016dxa ; He:2017vyf , which lack direct interpretation as distinguishability measures, or Schatten norms Basu:2017kzo ; He:2017txy , which are not effective measures of distinguishability in high-dimensional Hilbert spaces Zhang:2020mjv . Furthermore, fixed area state results depend on superposition details, obscuring universal quantum gravitational effects. Moreover, nonperturbative effects cannot explain observed $1/c$ corrections in holographic CFTs (one can see for example the footnote 17 in Kudler-Flam:2021alo ).
To resolve these issues, we compute directly the subsystem fidelity Nielsen:2010oan and quantum Jensen-Shannon divergence (QJSD) Majtey:2005ne ; Virosztek:2021fci between primary and thermal states in 2D holographic CFTs. Our methodology integrates three synergistic theoretical tools. Firstly, short-interval expansion leverages twist operators Calabrese:2004eu ; Cardy:2007mb ; Calabrese:2009qy and their operator product expansion (OPE) Headrick:2010zt ; Calabrese:2010he ; Rajabpour:2011pt ; Chen:2013kpa ; Ruggiero:2018hyl ; Lin:2016dxa ; Chen:2016lbu in two-dimensional conformal field theory to derive closed-form analytic expressions for subsystem fidelity and QJSD, establishing their quantitative roles in microstate distinguishability. Secondly, trace distance inequalities provide upper and lower bounds for $D(\rho_{A},\sigma_{A})$ through fidelity $F(\rho_{A},\sigma_{A})$ and QJSD $J(\rho_{A},\sigma_{A})$, bypassing computationally intractable trace norm calculations while preserving operational interpretability in high-dimensional Hilbert spaces. Thirdly, large $c$ expansions isolate $1/c$ corrections by prioritizing vacuum module dominance, thereby disentangling quantum gravitational signatures from universal classical parts in holographic systems.
For primary states $|\phi\rangle$, $|\psi\rangle$ with conformal weight difference $h_{\phi}-h_{\psi}=O(c^{0})$, we obtain the bound for the subsystem trace distance
$O(\ell^{4}/c)\lesssim D(\rho_{A,\phi},\rho_{A,\psi})\lesssim O(\ell^{2}/c^{1/2}),$ (1)
where $\ell$ is the size of the subsystem $A$. The large $c$ scaling breaks semiclassical indistinguishability at perturbative $1/c$ level. For a primary state $|\phi\rangle$ and thermal state $\rho_{\beta}$ with matching energy, we get
$O(c^{0}\ell^{8})\lesssim D(\rho_{A,\phi},\rho_{A,\beta})\lesssim O(c^{0}\ell^{4}).$ (2)
demonstrating quantum distinguishability despite identical classical geometries. Physically, our results imply that quantum gravity encodes sufficient information in small Hawking radiation subsystems to distinguish microstates. These findings directly challenge ETH for holographic systems Deutsch:2018thx ; Dymarsky:2018lhf , necessitating generalized Gibbs ensembles (GGEs) Rigol:2006jrd ; Cardy:2015xaa with Korteweg-de Vries (KdV) charges Bazhanov:1994ft ; Cardy:2015xaa ; Maloney:2018hdg , a framework supported by identity block operators Dymarsky:2019etq and large subsystem Rényi entropy Chen:2024lji ; Chen:2024ysb .
II Quantum state distinguishability measures
The trace distance between two quantum states with density matrices $\rho$ and $\sigma$ is Nielsen:2010oan
$D(\rho,\sigma)=\frac{1}{2}\textrm{tr}|\rho-\sigma|.$ (3)
By definition, $0\leq D(\rho,\sigma)\leq 1$, where $D=0$ implies identical states and $D=1$ implies orthogonal states. The trace distance quantifies the distinguishability of two states under optimal measurements.
The fidelity between $\rho$ and $\sigma$ is defined as Nielsen:2010oan
$F(\rho,\sigma)=\textrm{tr}\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}}.$ (4)
By definition, $0\leq F(\rho,\sigma)\leq 1$, where $F=1$ implies identical states and $F=0$ implies orthogonal states. Fidelity measures the closeness of two quantum states.
The trace distance and fidelity are equivalent metrics, linked by the Fuchs-van de Graaf inequalities Nielsen:2010oan
$1-F(\rho,\sigma)\leq D(\rho,\sigma)\leq\sqrt{1-F(\rho,\sigma)^{2}}.$ (5)
These inequalities show that trace distance and fidelity provide complementary bounds on the distinguishability of quantum states.
The relative entropy (Kullback-Leibler divergence) between $\rho$ and $\sigma$ is
$S(\rho,\sigma)=\textrm{tr}(\rho\log\rho)-\textrm{tr}(\rho\log\sigma).$ (6)
By definition, $S(\rho,\sigma)\geq 0$, and it is only well-defined when the support of $\rho$ is contained within the support of $\sigma$ (otherwise, $S(\rho,\sigma)=+\infty$).
The Quantum Jensen-Shannon Divergence (QJSD) Majtey:2005ne is a symmetric, bounded and smooth measure defined as
$J(\rho,\sigma)=S\Big{(}\frac{\rho+\sigma}{2}\Big{)}-\frac{1}{2}S(\rho)-\frac{1}{2}S(\sigma),$ (7)
where $S(\rho)=-\textrm{tr}(\rho\log\rho)$ is the von Neumann entropy. The QJSD can also be written as
$J(\rho,\sigma)=\frac{1}{2}\Big{[}S\Big{(}\rho\Big{\|}\frac{\rho+\sigma}{2}\Big{)}+S\Big{(}\sigma\Big{\|}\frac{\rho+\sigma}{2}\Big{)}\Big{]}.$ (8)
By definition, $0\leq J(\rho,\sigma)\leq\log 2$. The square root of QJSD is a rigorous metric for general mixed quantum states Lamberti:2008ptb ; Virosztek:2021fci ; Sra:2019twg .
The QJSD and trace distance are equivalent metrics, linked by the following inequalities
$\frac{J(\rho,\sigma)}{\log 2}\leq D(\rho,\sigma)\leq\sqrt{2J(\rho,\sigma)}.$ (9)
These bounds arise from Pinker’s inequality Watrous:2018rgz $D(\rho,\sigma)\leq\sqrt{\frac{1}{2}S(\rho\|\sigma)}$ and a result in Briet:2008lcr $J(\rho,\sigma)\leq(\log 2)D(\rho,\sigma)$.
III Subsystem distinguishability in 2D CFTs
Consider a subsystem $A$ of length $\ell$ in a two-dimensional (2D) conformal field theory (CFT) with total spatial length $L$. Let $\rho$ and $\sigma$ be two quantum states, with corresponding reduced density matrices (RDMs) $\rho_{A}$ and $\sigma_{A}$ obtained by tracing out the complement of subsystem $A$.
By employing twist operators Calabrese:2004eu ; Cardy:2007mb ; Calabrese:2009qy and their operator product expansion (OPE) Headrick:2010zt ; Calabrese:2010he ; Rajabpour:2011pt ; Chen:2013kpa ; Ruggiero:2018hyl ; Lin:2016dxa ; Chen:2016lbu , we derive the short interval expansion of the fidelity work
$F(\rho_{A},\sigma_{A})=1-\sum_{\mathcal{X}}\Big{[}\frac{(2\Delta_{\mathcal{X}})!\ell^{2\Delta_{X}}}{2^{4\Delta_{\mathcal{X}}+3}(\Delta_{\mathcal{X}}!)^{2}}\frac{(\langle\mathcal{X}\rangle_{\rho}-\langle\mathcal{X}\rangle_{\sigma})^{2}}{\mathrm{i}^{2s_{\mathcal{X}}}\alpha_{\mathcal{X}}}$
$+o(\ell^{2\Delta_{\mathcal{X}}})\Big{]},$ (10)
where the sum runs over all quasiprimary operators $\mathcal{X}$ characterized by conformal weights $(h_{\mathcal{X}},\bar{h}_{\mathcal{X}})$ and scaling dimension $\Delta_{\mathcal{X}}=h_{\mathcal{X}}+\bar{h}_{\mathcal{X}}$ and spin $s_{\mathcal{X}}=h_{\mathcal{X}}-\bar{h}_{\mathcal{X}}$. The coefficient $\alpha_{\mathcal{X}}$ is related to the normalization of the operator $\mathcal{X}$, and on the complex plane $\mathbb{C}$ the two-point function is given by
$\langle\mathcal{X}(z_{1},\bar{z}_{1})\mathcal{X}(z_{2},\bar{z}_{2})\rangle_{\mathbb{C}}=\frac{\alpha_{\mathcal{X}}}{(z_{1}-z_{2})^{2h_{\mathcal{X}}}(\bar{z}_{1}-\bar{z}_{2})^{2\bar{h}_{\mathcal{X}}}}.$ (11)
The fidelity (III) is consistent with the analytical and numerical results in free boson and fermion theories and XX chain and critical Ising chain Zhang:2019wqo ; Zhang:2019itb ; Zhang:2022nuh . Furthermore, in He:2017txy the subsystem QJSD was derived as
$J(\rho_{A},\sigma_{A})=\sum_{\mathcal{X}}\Big{[}\frac{\sqrt{\pi}\Gamma(\Delta_{\mathcal{X}}+1)\ell^{2\Delta_{X}}}{2^{2(\Delta_{\mathcal{X}}+2)}\Gamma(\Delta_{\mathcal{X}}+\frac{3}{2})}\frac{(\langle\mathcal{X}\rangle_{\rho}-\langle\mathcal{X}\rangle_{\sigma})^{2}}{\mathrm{i}^{2s_{\mathcal{X}}}\alpha_{\mathcal{X}}}$
$+o(\ell^{2\Delta_{\mathcal{X}}})\Big{]},$ (12)
with $\Gamma(x)$ denoting the Gamma function.
We choose $\mathcal{X}$ being one of the quasiprimary operators $\mathcal{X}$ satisfying $\langle\mathcal{X}\rangle_{\rho}\neq\langle\mathcal{X}\rangle_{\sigma}$ with the smallest scaling dimension $\Delta_{\mathcal{X}}$, and suppose that at the scaling dimension there is no other quasiprimary operators $\mathcal{Y}$ satisfying $\langle\mathcal{Y}\rangle_{\rho}\neq\langle\mathcal{Y}\rangle_{\sigma}$. Under such condition, the subsystem trace distance is Zhang:2019wqo ; Zhang:2019itb
$D(\rho_{A},\sigma_{A})=\frac{x_{\mathcal{X}}\ell^{\Delta_{\mathcal{X}}}}{2}\Big{|}\frac{\langle\mathcal{X}\rangle_{\rho}-\langle\mathcal{X}\rangle_{\sigma}}{\sqrt{\alpha_{\mathcal{X}}}}\Big{|}+o(\ell^{\Delta_{\mathcal{X}}}),$ (13)
with the coefficient
$x_{\mathcal{X}}=\lim_{p\to\frac{1}{2}}\frac{\mathrm{i}^{2ps_{\mathcal{X}}}}{\alpha_{\mathcal{X}}^{p}}\Big{\langle}\prod_{j=0}^{2p-1}\big{[}f_{j}^{h_{\mathcal{X}}}\bar{f}_{j}^{\bar{h}_{\mathcal{X}}}\mathcal{X}(f_{j},\bar{f}_{j})\big{]}\Big{\rangle}_{\mathbb{C}},~{}f_{j}\equiv\mathrm{e}^{\frac{\pi\mathrm{i}j}{p}}.$ (14)
In cases where the scaling dimension $\Delta_{\mathcal{X}}$ is degenerate, the form of the trace distance needs to be adjusted, although the scaling behavior $\ell^{\Delta_{\mathcal{X}}}$ remains unchanged.
The coefficient $x_{\mathcal{X}}$ is generally difficult to calculate. However, there are universal upper bounds of of the coefficient $z_{\mathcal{X}}$ from inequalities (5) and (9)
$x_{\mathcal{X}}\leq\frac{2^{-2\Delta}\sqrt{(2\Delta)!}}{\Delta!},$ (15)
$x_{\mathcal{X}}\leq\sqrt{\frac{\sqrt{\pi}G(\Delta+1)}{2^{2\Delta+1}G(\Delta+\frac{3}{2})}},$ (16)
with $G(x)$ denoting the Gamma function. The constraint from the fidelity (15) is tighter than the constraint from the fidelity (16), while the latter is the same as the constraint in Zhang:2019itb obtained from the relative entropy. Besides, in large $c$ limit, we can show that
$x_{T}\sim O(c^{0}),~{}x_{\mathcal{A}}\sim O(c^{0}),$ (17)
where $T$ is the holomorphic stress tensor and $\mathcal{A}=(TT)-\frac{3}{10}\partial^{2}T$ is a quasiprimay operator at level 4.
The expressions of the subsystem fidelity (III), QJSD (III) and trace distance (13) provide a systematic framework for quantifying the differences between the RDMs $\rho_{A}$ and $\sigma_{A}$ of subsystem $A$, capturing the distinguishability of black hole microstates.
IV Two black hole microstates
The 2D holographic CFT with a large central charge $c=\frac{3R}{2G}$ Brown:1986nw is dual to AdS3 quantum gravity, characterized by Newton’s constant $G$ and the AdS radius $R$. In this context, a primary state $|\phi\rangle$ with large conformal weights $(h_{\phi},\bar{h}_{\phi})$, where $h_{\phi}=\bar{h}_{\phi}=c\epsilon_{\phi}$ and $\epsilon_{\phi}>\frac{1}{24}$, corresponds to a microstate of a non-rotating BTZ black hole. Given that the holomorphic and anti-holomorphic sectors decouple, the subsequent analysis will focus solely on the contributions from the holomorphic sector.
We consider two primary states $|\phi\rangle$ and $|\psi\rangle$ with $h_{\phi}=c\epsilon_{\phi}$ and $h_{\psi}=c\epsilon_{\psi}$, where $\epsilon_{\phi}\neq\epsilon_{\psi}$. The contributions from the holomorphic vacuum conformal family to the subsystem fidelity and Quantum Jensen-Shannon Divergence (QJSD) are
$F(\rho_{A,\phi},\rho_{A,\psi})=1-\frac{3\pi^{4}c\ell^{4}(\epsilon_{\phi}-\epsilon_{\psi})^{2}}{32L^{4}}+o(\ell^{4}),$ (18)
$J(\rho_{A,\phi},\rho_{A,\psi})=\frac{2\pi^{4}c\ell^{4}(\epsilon_{\phi}-\epsilon_{\psi})^{2}}{15L^{4}}+o(\ell^{4}).$ (19)
For finite values of $\epsilon_{\phi}$, $\epsilon_{\psi}$ and $\epsilon_{\phi}-\epsilon_{\psi}\sim O(c^{0})$ in large $c$ limit, we have
$1-F(\rho_{A,\phi},\rho_{A,\psi})\sim O(c\ell^{4}),$ (20)
$J(\rho_{A,\phi},\rho_{A,\psi})\sim O(c\ell^{4}).$ (21)
According to the inequalities (5) and (9), the subsystem trace distance is bounded as
$O(c\ell^{4})\lesssim D(\rho_{A,\phi},\rho_{A,\psi})\lesssim O(c^{1/2}\ell^{2}).$ (22)
This result is intuitive, as two black holes with different masses have distinct metrics and are classically distinguishable. Although it is challenging to evaluate $D(\rho_{A,\phi},\rho_{A,\psi})$ explicitly, given (17) the expression for the subsystem trace distance (13) suggests
$D(\rho_{A,\phi},\rho_{A,\psi})=O(c^{1/2}\ell^{2}),$ (23)
which is consistent with the bounds (22).
We then consider two primary states $|\phi\rangle$ and $|\psi\rangle$ with conformal weights $h_{\phi}=c\epsilon_{\phi}+\delta_{\phi}$ and $h_{\psi}=c\epsilon_{\psi}+\delta_{\psi}$. For finite values of $\delta_{\phi}$, $\delta_{\psi}$ and $\delta_{\phi}-\delta_{\psi}\sim O(c^{0})$ in the large $c$ limit, the subsystem fidelity and QSJD are given by
$F(\rho_{A,\phi},\rho_{A,\psi})=1-\frac{3\pi^{4}\ell^{4}(\delta_{\phi}-\delta_{\psi})^{2}}{32cL^{4}}+o(\ell^{4}),$ (24)
$J(\rho_{A,\phi},\rho_{A,\psi})=\frac{2\pi^{4}\ell^{4}(\delta_{\phi}-\delta_{\psi})^{2}}{15cL^{4}}+o(\ell^{4}).$ (25)
Accordingly, the subsystem trace distance is bounded as
$O(\ell^{4}/c)\lesssim D(\rho_{A,\phi},\rho_{A,\psi})\lesssim O(\ell^{2}/c^{1/2}).$ (26)
The formula (13) indicates that
$D(\rho_{A,\phi},\rho_{A,\psi})=O(\ell^{2}/c^{1/2}).$ (27)
In this case, the two black hole microstates are classically indistinguishable, but they can be distinguished through perturbative quantum corrections. This result highlights the subtle quantum effects that allow for the differentiation of black hole microstates, even when classical metrics are identical.
For two primary states $|\phi\rangle$ and $|\psi\rangle$ with the same conformal weights $h_{\phi}=h_{\psi}$ and $\bar{h}_{\phi}=\bar{h}_{\psi}$, the vacuum conformal family do not contribute to the short interval expansion of the subsystem fidelity and QJSD. With $\phi$ and $\psi$ being hermitian and the normalization $\alpha_{\phi}=\alpha_{\psi}=1$, we obtain
$F(\rho_{A,\phi},\rho_{A,\psi})=1-\sum_{\mathcal{X}}\Big{[}\frac{(2\Delta_{\mathcal{X}})!}{2^{2\Delta_{\mathcal{X}}+3}(\Delta_{\mathcal{X}}!)^{2}}\Big{(}\frac{\pi\ell}{L}\Big{)}^{2\Delta_{\mathcal{X}}}$
×$\times\frac{(C_{\phi\mathcal{X}\phi}-C_{\psi\mathcal{X}\psi})^{2}}{\alpha_{\mathcal{X}}}+o(\ell^{2\Delta_{\mathcal{X}}})\Big{]},$ (28)
$J(\rho_{A,\phi},\rho_{A,\psi})=\sum_{\mathcal{X}}\Big{[}\frac{\sqrt{\pi}\Gamma(\Delta_{\mathcal{X}}+1)}{16\Gamma(\Delta_{\mathcal{X}}+\frac{3}{2})}\Big{(}\frac{\pi\ell}{L}\Big{)}^{2\Delta_{\mathcal{X}}}$
×$\times\frac{(C_{\phi\mathcal{X}\phi}-C_{\psi\mathcal{X}\psi})^{2}}{\alpha_{\mathcal{X}}}+o(\ell^{2\Delta_{\mathcal{X}}})\Big{]},$ (29)
with $C_{\phi\mathcal{X}\phi}$ and $C_{\psi\mathcal{X}\psi}$ being the structure constants. Note that $1-F(\rho_{A,\phi},\rho_{A,\psi})$ and $J(\rho_{A,\phi},\rho_{A,\psi})$ are at the same leading order, whose behavior in the large $c$ limit depends on the details of the structure constants in the 2D CFT.
V Black hole microstate and thermal state
Finally, we examine a primary state $|\phi\rangle$ with conformal weight $h_{\phi}=c\epsilon_{\phi}$ and a thermal state $\rho_{\beta}=\mathrm{e}^{-\beta H}/Z(\beta)$. According to Fitzpatrick:2014vua ; Fitzpatrick:2015zha , when
$\beta=\frac{L}{\sqrt{24\epsilon_{\phi}-1}},$ (30)
the two states have the same energy and the same order $c$ entanglement entropy. However, He:2017vyf shows that with $1/c$ corrections, their entanglement entropies differ, which also indicates that the subsystem relative entropy $S(\rho_{A,\phi}\|\rho_{A,\beta})$ scales as $O(c^{0}\ell^{8})$.
We obtain the subsystem fidelity as
$F(\rho_{A,\phi},\rho_{A,\beta})=1-\frac{7\pi^{8}c\ell^{8}(22\epsilon_{\phi}-1)^{2}\epsilon_{\phi}^{2}}{512(5c+22)L^{8}}+o(\ell^{8}).$ (31)
The subsystem Quantum Jensen-Shannon Divergence (QJSD) was derived in He:2017txy as
$J(\rho_{A,\phi},\rho_{A,\beta})=\frac{32\pi^{8}c\ell^{8}\epsilon_{\phi}^{2}(22\epsilon_{\phi}-1)^{2}}{1575(5c+22)L^{8}}+o(\ell^{8}).$ (32)
Applying the inequalities (5) and (9), we find the subsystem trace distance is bounded by
$O(c^{0}\ell^{8})\lesssim D(\rho_{A,\phi},\rho_{A,\beta})\lesssim O(c^{0}\ell^{4}).$ (33)
The formula for the subsystem trace distance (13) indicates
$D(\rho_{A,\phi},\rho_{A,\beta})=O(c^{0}\ell^{4}).$ (34)
These results directly demonstrate that the black hole microstate and the canonical thermal state are perturbatively distinguishable. This perturbative distinguishability highlights the subtle quantum differences between the two states, even though they may appear similar at classical level.
VI Discussions and Conclusions
Our results from holographic CFTs establish that quantum gravitational corrections can induce perturbative distinguishability between black hole microstates and thermal ensembles, even for small subsystems. This conclusion rests on three pillars: 1. Subsystem fidelity and trace distance as operational metrics: By deriving the short-interval expansions of fidelity $F(\rho_{A},\sigma_{A})$ and QJSD $J(\rho_{A},\sigma_{A})$, we bypass the limitations of entropy-based measures, directly quantifying state distinguishability through rigorous inequalities for the trace distance. 2. Primary states as exact microstates: Unlike fixed area states, superpositions with ambiguous physicality, primary states represent exact eigenstates, isolating intrinsic quantum gravitational effects. 3. Universal $1/c$ scaling: The scaling of $D(\rho_{A},\sigma_{A})$ demonstrate that distinguishability persists in for a finite central charge, provided $\ell/L$ scales appropriately.
The failure of traditional ETH for primary states, evident from the subsystem trace distances between microstates and thermal ensembles, necessitates extending thermalization to generalized Gibbs ensembles (GGEs) incorporating KdV charges. This aligns with recent findings in large-$c$ CFTs Dymarsky:2019etq ; Chen:2024lji ; Chen:2024ysb , where KdV conservation laws govern subsystem thermalization.
The perturbative distinguishability of microstates via small subsystems resolves a key obstacle to information retrieval: while semiclassical arguments forbid distinguishing microstates from Hawking radiation, quantum corrections at $1/c$ order encode sufficient information in finite-dimensional subsystems. This mechanism complements the island formula by providing an operational pathway to reconstruct microstate details from radiation.
While primary states dominate the microcanonical ensemble, descendants Guo:2018pvi ; Guo:2018fnv ; Brehm:2020zri may exhibit enhanced distinguishability due to Virasoro hair. Extending our framework to include descendant contributions could reveal new distinguishability structures. While our analysis focuses on 2D CFTs, analogous results for higher-dimensional black holes could test the universality of perturbative distinguishability.
In summary, this work demonstrates that quantum gravity corrections in holographic CFTs fundamentally alter the semiclassical notion of black hole thermalization. By establishing perturbative distinguishability as a generic feature of AdS3/CFT2, we bridge the gap between information-theoretic principles and geometric unitarity, offering a concrete step toward fully resolving the black hole information paradox.
Acknowledgements
The author acknowledgements support from the National Natural Science Foundation of China (NSFC) grant number 12205217.